OpenStudy (anonymous):

Solve the equation 1 over 3 x + 2 = 10 for x. 36 30 24 12

OpenStudy (anonymous):

how do i do this?

OpenStudy (anonymous):

You should be able to move the 10 over and set the equation equal to 0. Once you do that change 10 into a fraction. You should end up with (1/(3x+2)) - 10/1 Now make it so it has a common denominator by multiplying it by 3x+2. Then solve.

OpenStudy (anonymous):

lol im confused now

OpenStudy (anonymous):

oh ok so then X would equal 28?

OpenStudy (anonymous):

\[\frac{ 1 }{ 3x+2 }= 10\] \[\frac{ 1 }{3x+2 }-\frac{ 10 }{ 1 }=0\] \[\frac{ 1 }{ 3x+2 }-(\frac{ 10 }{ 1 }\times \frac{ 3x+2 }{3x+2 })\]

OpenStudy (anonymous):

so x= 28?

OpenStudy (anonymous):

Is that one of your answer options?

OpenStudy (anonymous):

oops nope lol

OpenStudy (anonymous):

HAHAHAHA I suck at this

OpenStudy (anonymous):

Apparently I do too, I keep getting an improper fraction :( Hold on, let me try it a different way

OpenStudy (anonymous):

I multiplied by the resiprical then subtracted 2 on both sides

OpenStudy (anonymous):

Does only the 3x+2 = 10 or does the entire equation = 10?

OpenStudy (anonymous):

so it would go from this : 1 over 3 x + 2 = 10 for x. to this: x+2=30 to this: x=28 but apparently its not right

OpenStudy (anonymous):

Well you dropped an x I think. It should be 30x

OpenStudy (anonymous):

no I left the x where it was but that may have been my mistake

OpenStudy (anonymous):

no then it would be x= 0.0666666666666666666666

OpenStudy (anonymous):